OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = (2 points)

OpenStudy (anonymous):

@Luigi0210 @agent0smith

OpenStudy (anonymous):

@alassen

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

this is going to take few steps

ganeshie8 (ganeshie8):

f(n) = f(1) + f(2) + f(n - 1)

OpenStudy (anonymous):

O crap lmao

ganeshie8 (ganeshie8):

you want to find f(5), so plugin n = 5 above ^

ganeshie8 (ganeshie8):

yes :)

ganeshie8 (ganeshie8):

f(n) = f(1) + f(2) + f(n - 1) f(5) = ?

ganeshie8 (ganeshie8):

f(5) = f(1) + f(2) + f(5-1) = 2 + 4 + f(4) = 6 + f(4)

ganeshie8 (ganeshie8):

do we knw f(4) yet ?

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

O my gosh don't think im stupid

ganeshie8 (ganeshie8):

lets find out

ganeshie8 (ganeshie8):

but before diving in to find f(4), just save wat we have till now, so that we can come back here once we find f(4) :- f(5) = 6 + f(4) ------------------------------------- (1)

ganeshie8 (ganeshie8):

we wil get back to this equation (1), after finding f(4), okay ?

OpenStudy (anonymous):

ummm okay

ganeshie8 (ganeshie8):

f(n) = f(1) + f(2) + f(n - 1) we want to find f(4), so we plugin n = 4 :- f(4) = ?

OpenStudy (anonymous):

16?

ganeshie8 (ganeshie8):

f(4) = f(1) + f(2) + f(4-1) = 2 + 4 + f(3) = 6 + f(3)

ganeshie8 (ganeshie8):

we're not that lucky yet ;)

ganeshie8 (ganeshie8):

so, we got f(4) = 6 + f(3) --------------------------------(2)

ganeshie8 (ganeshie8):

since we dont knw f(3) yet, we need to find f(3) value first, and get back to this equation, okay ?

ganeshie8 (ganeshie8):

u still wid me ha

OpenStudy (anonymous):

Yeah I think so......

ganeshie8 (ganeshie8):

good :) lets find f(3) :- f(n) = f(1) + f(2) + f(n - 1) plugin n = 3 f(3) = f(1) + f(2) + f(3-1) = 2 + 4 + f(2) = 2 + 4 + 4 = 10

ganeshie8 (ganeshie8):

looks we're done. we got f(3) = 10, so we can go back to 2nd equation and plug this f(3) = 10 value

OpenStudy (anonymous):

What there's more O boy!!!

ganeshie8 (ganeshie8):

remember we saved two equations before

ganeshie8 (ganeshie8):

go back to 2nd equation, and plugin f(3) = 10

ganeshie8 (ganeshie8):

that gives u f(4) value

ganeshie8 (ganeshie8):

f(4) = 6 + f(3) --------------------------------(2) f(4) = 6 + 10 = 16

ganeshie8 (ganeshie8):

plug this f(4) = 16 in equation (1)

ganeshie8 (ganeshie8):

f(5) = 6 + f(4) ------------------------------------- (1)

ganeshie8 (ganeshie8):

f(5) = 6 + 16 = 22

ganeshie8 (ganeshie8):

we're done.

OpenStudy (anonymous):

Wow I actually kind of got all that toward the end thank you I wish I knew all that so I wouldn't need help lol thanks again :)

ganeshie8 (ganeshie8):

glad to hear that :) im sure if u do one more problem like this u wil feel more confident... :)

OpenStudy (anonymous):

:) thank you

ganeshie8 (ganeshie8):

np.. u wlc :)